Find a counterexample to show that the following statement is false:
The sum of two five-digit numbers is a six-digit number

Answer by KMST(5328)

Some pairs of five-digit numbers add up to a six-digit number.
However, there are many pairs of five-digit numbers that add up to a five-digit number, and any of those pairs could be your counterexample.
10,000+10,000=20,000
10,000+20,000=30,000
12,345+67,890=80,235